Fermi National Laboratory
 

Operation Summary for June 20, 2011 - June 22, 2011

 

Monday June 20

The day shift began with Operations monitoring the Tevatron (TEV) store 8827 (98.01E30) and the Recycler stash (290.3E10), with the Antiproton Source (Pbar) stacking (8.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, NuMI personnel began 40 minutes of target scans at 11:07 AM. Operations terminated store 8827 at 1:09 PM, put the TEV into shot setup, and established store 8828 at 2:49 PM with an average initial luminosity (AIL) of 392.65E30. Operators reported that the proton load for the store was delayed due to the need of cleaning the TEV sector B4 pond water pump strainer. Linac personnel began conducting a 20-minute destructive beam study at 2:49 PM.

During the evening shift, Operations disabled beam to T-978 at 5:10 PM. At 6:30 PM, the Booster beam permit tripped off with no indication as to why. Operators investigated and found the rack containing the Booster RF station (BRF15) equipment had no power. RF technicians came in and had power temporarily restored to the rack by 8:18 PM. Operators reported that a TEV injection kicker tripped off on a filament fault. Operators contacted a kicker person who told them to reset the fault.

Tuesday June 21

The midnight shift began with Operations monitoring store 8828 (109.28E30) and the stash (210.2E10), with Pbar stacking (13.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:30 AM, operators reported that the Booster West Anode power supply tripped off for the first of seven times in four hours; it reset okay. Operations turned off the NuMI beam at 3:30 AM in preparation for an access during the day shift. Operations established beam to T-978 at 5:01 AM.

During the day shift, Operations terminated store 8828 at 8:42 AM, put the TEV into shot setup, and established store 8829 at 10:02 AM with an AIL of 386.34E30. Operators prepared the Pelletron for an access at 10:25 AM for safety system tests; Interlock technicians completed their tests at 11:19 AM. At 12:38 PM, a Booster RF station (BRF16) tripped off for the first of four times in 20 minutes. Operators bypassed the station so that RF personnel could investigate. NuMI personnel began target scans and MTA - MuCool began taking beam. Interlock techs began Neutron Therapy Facility safety system checks at 2:08 PM; they completed the test at 2:18 PM. RF personnel reported at 2:37 PM that they had replaced a power supply for BRF16 and brought the station back online. Pbar personnel also repaired an Accumulator RF station (ARF1) during the Booster downtime. TEV personnel began a luminosity leveling study at 2:43 PM. At 3:02 PM, operators reported a power glitch that took down the MI RF system and the Low-Level RF system; operators rebooted the systems.

During the evening shift, Operations stopped sending beam to MTA at 4:38 PM. NuMI completed their target scans at 5:16 PM. Operations disabled beam to T-978 at 6:08 PM. TEV personnel completed their leveling studies at 6:31 PM. At 7:48 PM, the MiniBooNE Horn tripped off on a power supply fault twice in two minutes; it reset. Operators prepared NuMI for an access. At 8:05 PM, a tornado warning caused operators to turn off the Proton Source and go to the tornado shelter. The TEV remained operating, but CDF and D0 turned off their detectors. Operators came back to the Main Control Room and brought the Proton Source back on at 9:07 PM.

Wednesday June 22

The midnight shift began with Operations monitoring store 8829 (77.84E30) and the stash (298.7E10), with Pbar stacking (13.5E10), with NuMI off for the night, and with MiniBooNE taking beam.

During the midnight shift, at 3:08 AM, Operations lowered the beam intensity for MiniBooNE. Operations terminated store 8829 at 4:49 AM, put the TEV into shot setup, and established store 8830 at 6:13 AM with an AIL of 309.3E30.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a coalescing study. She will also allow CDF and D0 personnel to access their collision halls for two hours this evening after Operations terminates store 8830. NuMI personnel had reported some time ago that their target was leaking helium used to pressurize the channel that encloses and separates the target from its cooling system. They have found a small crack on their target tip, which is possibly where the helium is leaking from, but report that this will not stop NuMI from running. NuMI will resume taking beam later this morning.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/22/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory